# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 2.3. Special divisors
(2.9) $\left\langle x_{1},x_{2}\right\rangle=\iota_{0}^{-1}(\lambda_{0}^{-1}\circ x_{2}^{\vee}\circ\lambda\circ x_{1})\in\mathcal{O}_{\boldsymbol{k}}$
[…]
@@
    $x\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}\left(A_{0},A\right)$ satisfies $\langle x,x\rangle=m$.
According to the discussion in [${}^{+}$20, §2.5], we may regard $\mathcal{Z}(m)$ as a Cartier divisor on $\mathcal{S}$.
2.4. Toroidal compactification
In this subsection, we briefly review the theory of toroidal compatification of $\mathcal{S}$ following [${}^{+}$20, §3].
A proper cusp label is an isomorphism class of pairs $\Phi=(\mathfrak{n},L_{\Phi})$ in which $\mathfrak{n}$ is a projective $\mathcal{O}_{\boldsymbol{k}}$-module of rank one, and $L_{\Phi}$ is a unimodular Hermitian lattice of signature $(n-2,0)$. Consider a pair $\mathfrak{m}\subset L$ where $L$ is a unimodular hermitian lattice of signature $(n-1,1)$ and $\mathfrak{m}$ is an isotropic direct summand of rank one. A normal decomposition of $\mathfrak{m}\subset L$ is an $\mathcal{O}_{\boldsymbol{k}}$-module direct sum decomposition
(2.10) $L=(\mathfrak{m}\oplus\mathfrak{n})\obot L_{\Phi}$
where $L_{\Phi}=(\mathfrak{m}\oplus\mathfrak{n})^{\perp}$ and $\mathfrak{n}$ is an isotropic direct summand of rank one and can be identified as $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathfrak{m},\mathcal{O}_{\boldsymbol{k}})$. The hermitian form on $L$ makes $L_{\Phi}\cong$ $\mathfrak{m}^{\perp}/\mathfrak{m}$ into a unimodular hermitian lattice of signature $(n-2,0)$.
@@ 2.4. Toroidal compactification
[How15, Proposition 2.6.3] Every pair $\mathfrak{m}\subset L$ as above admits a normal decomposition. The rule
$\mathfrak{m}\subset L\mapsto\left(L/\mathfrak{m}^{\perp},\mathfrak{m}^{\perp}/\mathfrak{m}\right)\cong(\mathfrak{n},L_{\Phi})$
establishes a bijection between the isomorphism classes of pairs $\mathfrak{m}\subset L$ as above, and the set of cusp labels.
Now let $L$ be the unimodular lattice defined in 2.5.
@@
We define $\mathrm{Cusp}(L)$ to be the set of proper cusp labels $\Phi=(\mathfrak{n},L_{\Phi})$ such that there exists a $\mathfrak{m}$ with $L=\mathfrak{m}\oplus\mathfrak{n}\obot L_{\Phi}$. Moreover, we denote $\mathrm{Cusp}([[L]])=\coprod_{L^{\prime}\in[[L]]}\mathrm{Cusp}(L^{\prime})$.
@@
Because of Lemma 2.3, we also denote a proper cusp label as $\Phi=(\mathfrak{m},L)$. The definition of proper cusp labels in [${}^{+}$20] is different with Definition 2.4. However, according to [${}^{+}$20, Lemma 3.1.4] and Lemma 2.3, there is a natural bijection between the set of equivalence classes of proper cusp labels defined in [${}^{+}$20] and the one defined in Definition 2.4.
We describe the boundary more explicitly following [${}^{+}$20, §3.3]. For $\left(A_{0},B,\ldots\right)\in\mathcal{M}_{(1,0)}(S)\times_{\mathcal{O}_{\boldsymbol{k}}}\mathcal{M}_{(n-2,0)}(S)$, the étale sheaf $\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}\left(B,A_{0}\right)$ is locally constant by [BHY15, Theorem 5.1]. For a fixed cusp label $\Phi=(\mathfrak{n},L_{\Phi})$, let $\Lambda_{\Phi}=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(L_{\Phi},\mathcal{O}_{\boldsymbol{k}})$ be the hermitian dual of $L_{\Phi}$. Let $\mathcal{A}_{\Phi}$ be the moduli space of triples $\left(A_{0},B,\varrho\right)$ over $\mathcal{O}_{\boldsymbol{k}}$-schemes $S$, where
@@
$\left(A_{0},B,\ldots\right)\in\mathcal{M}_{(1,0)}(S)\times_{\mathcal{O}_{\boldsymbol{k}}}\mathcal{M}_{(n-2,0)}(S),$
@@
is an isomorphism of étale sheaves of hermitian $\mathcal{O}_{\boldsymbol{k}}$-modules. Then $\mathcal{A}_{\Phi}\rightarrow\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$ is smooth of relative dimension 00. Now we define $\mathcal{B}_{\Phi}$ to be the moduli space of quadruples $\left(A_{0},B,\varrho,c\right)$ over the $\mathcal{O}_{\boldsymbol{k}}$-schemes $S$, where $\left(A_{0},B,\varrho\right)\in\mathcal{A}_{\Phi}(S)$ and $c:\mathfrak{n}\rightarrow B$ is an $\mathcal{O}_{\boldsymbol{k}}$-linear homomorphism of group schemes over $S$. In other words,
where $\left(A_{0},B,\varrho\right)$ is the universal object over $\mathcal{A}_{\Phi}$. More explicitly, according to [${}^{+}$20, Proposition 3.4.4], we have
(2.12) $\mathcal{B}_{\Phi}\cong A_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}L_{\Phi},$
where $A_{0}$ is the universal object of $\mathcal{M}_{(1,0)}$ and $\otimes$ is Serre’s tensor construction. The forgetful morphism $\mathcal{B}_{\Phi}\rightarrow\mathcal{A}_{\Phi}$ is smooth of relative dimension $n-2$.
Suppose we fix $\mu,\nu\in\mathfrak{n}$. For any scheme $U$ and any morphism $U\rightarrow\mathcal{B}_{\Phi}$, there is a corresponding quadruple $(A_{0},B,\varrho,c)$ over $U$. Evaluating the morphism of $U$-group schemes $c:\mathfrak{n}\rightarrow B$ at $\mu$ and $\nu$ determines $U$-points $c(\mu),c(\nu)\in B(U)$, and hence determines a morphism of $U$-schemes
$U\xrightarrow[]{c(\mu)\times c(\nu)}B\times B\cong B\times B^{\vee}.$
Denote by $\mathcal{L}(\mu,\nu)_{U}$ the pullback of the Poincaré line bundle under this morphism. As $U$ varies, these line bundles are obtained as the pullback of a single line bundle $\mathcal{L}(\mu,\nu)$ over $\mathcal{B}_{\Phi}$.
It follows from standard bilinear properties of Poincaré bundles and $\mathcal{O}_{\boldsymbol{k}}$-linearality of the polarization of $B$ that $\mathcal{L}(\mu,\nu)$, up to canonical isomorphism, only depends on the image of $\mu,\nu$ in
@@
Our assumption that $D$ is odd implies that $\mathrm{Sym}_{\Phi}$ is a free $\mathbb{Z}$-module of rank $1$. Moreover there is a positive cone in $\mathrm{Sym}_{\Phi}\otimes_{\mathbb{Z}}\mathbb{R}$ uniquely determined by the condition $\mu\otimes\mu\geq 0$ for all $\mu\in\mathfrak{n}$. Thus all the line bundles $\mathcal{L}(\chi)$ are powers of the distinguished line bundle
$\mathcal{L}_{\Phi}=\mathcal{L}(\chi_{0}),$
determined by the unique positive generator $\chi_{0}\in\mathrm{Sym}_{\Phi}$.

### v2

@@ 2.3. Special divisors
(2.9) $\left(x_{1},x_{2}\right)=\iota_{0}^{-1}(\lambda_{0}^{-1}\circ x_{2}^{\vee}\circ\lambda\circ x_{1})\in\mathcal{O}_{\boldsymbol{k}}$
where $x_{2}^{\vee}:A^{\vee}\rightarrow A_{0}^{\vee}$ is the dual homomorphism of $x_{2}$. By [KR14, Lemma 2.7], the form $\left(,\right)$ is positive-definite.
@@
    $x\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}\left(A_{0},A\right)$ satisfies $(x,x)=m$.
According to the discussion in [BHK+20, §2.5], we may regard $\mathcal{Z}(m)$ as a Cartier divisor on $\mathcal{S}$.
In the rest of the section, we recall the construction of the toroidal compactification of $\mathcal{S}$ following [BHK+20, §3] and [How15, §2].
2.4. Cusp labels
A proper cusp label is an isomorphism class of pairs $\Phi=(\mathfrak{n},L_{\Phi})$ in which $\mathfrak{n}$ is a projective $\mathcal{O}_{\boldsymbol{k}}$-module of rank one, and $L_{\Phi}$ is a unimodular Hermitian lattice of signature (n2,0)(n-2,0). Consider a pair $\mathfrak{m}\subset M$ where $M$ is a unimodular hermitian lattice of signature (n1,1)(n-1,1) and $\mathfrak{m}$ is an isotropic direct summand of rank one. A normal decomposition of $\mathfrak{m}\subset M$ is an $\mathcal{O}_{\boldsymbol{k}}$-module direct sum decomposition
(2.10) $M=(\mathfrak{m}\oplus\mathfrak{n})\obot L_{\Phi}$
where $L_{\Phi}=(\mathfrak{m}\oplus\mathfrak{n})^{\perp}$ and $\mathfrak{n}$ is an isotropic direct summand of rank one and can be identified as $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathfrak{m},\mathcal{O}_{\boldsymbol{k}})$. The hermitian form on $M$ makes $L_{\Phi}\cong$ $\mathfrak{m}^{\perp}/\mathfrak{m}$ into a unimodular hermitian lattice of signature (n2,0)(n-2,0).
@@ 2.4. Cusp labels
[How15, Proposition 2.6.3] Every pair $\mathfrak{m}\subset M$ as above admits a normal decomposition. The rule
$\mathfrak{m}\subset M\mapsto\left(M/\mathfrak{m}^{\perp},\mathfrak{m}^{\perp}/\mathfrak{m}\right)\cong(\mathfrak{n},L_{\Phi})$
establishes a bijection between the isomorphism classes of pairs $\mathfrak{m}\subset M$ as above, and the set of cusp labels.
Recall that $L$ is the unimodular lattice defined in 2.5.
@@
We define $\mathrm{Cusp}(M)$ to be the set of proper cusp labels $\Phi=(\mathfrak{n},L_{\Phi})$ such that there exists a $\mathfrak{m}$ with $M=\mathfrak{m}\oplus\mathfrak{n}\obot L_{\Phi}$. Moreover, we denote $\mathrm{Cusp}([[L]])=\coprod_{M\in[[L]]}\mathrm{Cusp}(M)$.
@@
Because of Lemma 2.3, we also denote a proper cusp label as $\Phi=(\mathfrak{m}\subset M)$. The definition of proper cusp labels in [BHK+20] is different with Definition 2.4. However, according to [BHK+20, Lemma 3.1.4] and Lemma 2.3, there is a natural bijection between the set of equivalence classes of proper cusp labels defined in [BHK+20] and the one defined in Definition 2.4. Indeed, assume $(P,g)$ is as in [BHK+20, Definition 3.1.1], then $gL=M$ is a self-dual $\mathcal{O}_{\boldsymbol{k}}$-lattice and $P$ determines an isotropic line $J\subset W$. Then $\mathrm{Hom}_{\boldsymbol{k}}(W_{0},J)\cap L$ determines an isotropic dirrect summand of rank one. Then [BHK+20, Lemma 3.1.4] shows that this map is an injection. Now given $(\mathfrak{m},M)$, we can choose $g$ such that $gL=M$. Then we can choose $J\subset W$ such that $\mathrm{Hom}_{\boldsymbol{k}}(W_{0},J)\cap L=\mathfrak{m}$ and set $g=\mathrm{Stab}_{G}(J)$.
2.5. Degenerating abelian schemes
We review the theory of degenerating abelian scheme following [How15, §2.3] and [Lan13, §5.1]. For a projective $\mathcal{O}_{\boldsymbol{k}}$-module $\mathfrak{p}$ of rank $1$, let $\underline{\mathfrak{p}}$ be the associated constant $\mathcal{O}_{\boldsymbol{k}}$-module scheme over $\mathrm{Spec}\,\mathbb{Z}$. Let $X$ be an $\mathcal{O}_{\boldsymbol{k}}$-stack, $Z\rightarrow X$ be a closed substack, and $U\subset X\setminus Z$ be a dense open substack. We remark that we can also take $U$ to be the generic point of $X$ when $X$ is an irreducible scheme and the following discussion will be the same. Let $\Phi=(\mathfrak{n},L_{\Phi})$ or $(\mathfrak{m}\subset M)$ be a cusp label as in §2.4. Let $\Lambda_{\Phi}=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(L_{\Phi},\mathcal{O}_{\boldsymbol{k}})$ be the hermitian dual of $L_{\Phi}$.
Definition 2.5.
A semiabelian scheme over $X$ is a smooth commutative group scheme $G\rightarrow X$, such that for every geometric point $z\rightarrow X$ the fiber $G_{z}$ is an extension
$0\rightarrow T\rightarrow G_{z}\rightarrow B\rightarrow 0$
of an abelian variety by a torus.
Definition 2.6.
A degenerating abelian scheme of type $\Phi$ relative to $(X,Z,U)$ is a triple $(G,\iota,\lambda)$ such that
  • $G$ is a semi-abelian scheme over $X$ such that $G_{U}$ is an abelian scheme;
  • $\iota:\mathcal{O}_{\boldsymbol{k}}\rightarrow\operatorname{End}(G_{U})$ is an action of $\mathcal{O}_{\boldsymbol{k}}$ on $G_{U}$;
  • $\lambda:G_{U}\rightarrow G_{U}^{\vee}$ is a principal polarization whose induced Rosati involution $\dagger$ on $\operatorname{End}^{0}(G_{U})$ satisfies $\iota(\alpha)^{\dagger}=\iota(\bar{\alpha})$ for all $\alpha\in\mathcal{O}_{\boldsymbol{k}}$;
  • there is an abelian scheme $B_{Z}$ over $Z$ equipped with an $\mathcal{O}_{\boldsymbol{k}}$-action, and an $\mathcal{O}_{\boldsymbol{k}}$-linear exact sequence
    $0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathbb{G}_{m}\rightarrow G_{Z}\rightarrow B_{Z}\rightarrow 0.$
If in addition
  • there is an $\mathcal{O}_{\boldsymbol{k}}$-stable $\mathcal{O}_{U}$-module local direct summand $\mathcal{F}\subset\operatorname{Lie}(G_{U})$ of rank n1n-1 satisfying Krämer’s condition as in Condition (4) in the definition of $\mathcal{M}_{(r,s)}$;
  • $(A_{0},\iota_{0},\lambda_{0})\in\mathcal{M}_{(1,0)}(X)$;
  • there is an isomorphism of étale sheaves of hermitian $\mathcal{O}_{\boldsymbol{k}}$-modules over $Z$:
@@ 2.5. Degenerating abelian schemes
    $\underline{\Lambda_{\Phi}}\cong\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}\left(B_{Z},A_{0}\right);$
then we say $(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F})$ is a degenerating abelian scheme of type $\Phi$ and signature (n1,1)(n-1,1) relative to $(X,Z,U)$. We denote the category of degenerating abelian scheme of type $\Phi$ (and signature (n1,1)(n-1,1) resp.) relative to $(X,Z,U)$ as $\mathrm{DEG}^{\Phi}(X,Z,U)$ ($\mathrm{DEG}^{\Phi}_{(n-1,1)}(X,Z,U)$ resp.), with isomorphisms in the obvious sense being morphisms.
Definition 2.7.
Degeneration data of type $\Phi$ relative to $(X,Z,U)$ consist of tuples $(B,\kappa,\psi,c,c^{\vee},\tau)$ such that
  • $B\rightarrow X$ is an abelian scheme;
  • $\kappa:\mathcal{O}_{\boldsymbol{k}}\rightarrow\operatorname{End}(B)$ is an action of $\mathcal{O}_{\boldsymbol{k}}$ on $B$;
  • $\psi:B\rightarrow B^{\vee}$ is a principal polarization whose induced Rosati involution $\dagger$ on $\operatorname{End}^{0}(B)$ satisfies $\iota(\alpha)^{\dagger}=\iota(\bar{\alpha})$ for all $\alpha\in\mathcal{O}_{\boldsymbol{k}}$;
  • $c:\underline{\mathfrak{n}}_{/X}\rightarrow B^{\vee}$ and $c^{\vee}:\underline{\mathfrak{n}}_{/X}\rightarrow B$ are $\mathcal{O}_{\boldsymbol{k}}$-module maps satisfying $c=\psi\circ c^{\vee}$;
  • $\tau$ is a positive, symmetric, and $\mathcal{O}_{\boldsymbol{k}}$-linear isomorphism
    (2.11) $\tau:1_{(\underline{\mathfrak{n}}\times\underline{\mathfrak{n}})|_{U}}\rightarrow(c^{\vee}\times c)^{*}(\mathcal{P}^{-1})|_{(\underline{\mathfrak{n}}\times\underline{\mathfrak{n}})|_{U}}$
    of $\mathbb{G}_{m}$-biextensions of $(\underline{\mathfrak{n}}\times\underline{\mathfrak{n}})|_{U}$ (see below). Here $\mathcal{P}$ is the Poincaré sheaf on $B\times B^{\vee}$.
If in addition
  • $(A_{0},\iota_{0},\lambda_{0})\in\mathcal{M}_{(1,0)}(X)$;
  • $B\in\mathcal{M}_{(n-2,0)}(X)$ and there is an isomorphism of étale sheaves of hermitian $\mathcal{O}_{\boldsymbol{k}}$-modules over $X$:
    $\underline{\Lambda_{\Phi}}\cong\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}\left(B,A_{0}\right);$
then we say $(A_{0},\iota_{0},\lambda_{0},B,\kappa,\psi,c,c^{\vee},\tau)$ is degeneration data of type $\Phi$ and signature (n1,1)(n-1,1) relative to $(X,Z,U)$. We denote the category of degeneration data of type $\Phi$ (and signature (n1,1)(n-1,1) resp.) relative to $(X,Z,U)$ as $\mathrm{DD}^{\Phi}(X,Z,U)$ ($\mathrm{DD}^{\Phi}_{(n-1,1)}(X,Z,U)$ resp.), with isomorphisms in the obvious sense being morphisms.
We explain the meaning of $\tau$ in more detail. To give a $\mathbb{G}_{m}$-biextension on $\underline{\mathfrak{n}}\times\underline{\mathfrak{n}}|_{X}$ is equivalent to giving a collection of invertible sheaves $\mathcal{E}(\mu,\nu)_{(\mu,\nu)\in\mathfrak{n}\times\mathfrak{n}}$ on $X$, together with isomorphisms
$\mathcal{E}(\mu_{1}+\mu_{2},\nu)\cong\mathcal{E}(\mu_{1},\nu)\otimes\mathcal{E}(\mu_{2},\nu)$
@@
$\mathcal{E}(\mu,\nu_{1}+\nu_{2})\cong\mathcal{E}(\mu,\nu_{1})\otimes\mathcal{E}(\mu,\nu_{2})$
satisfying certain partial group axioms. Denote by $\mathcal{L}(\mu,\nu)$ the pullback of the Poincaré line bundle under the morphism $c^{\vee}(\mu)\times c(\nu):X\rightarrow B\times B^{\vee}$. It follows from the standard bilinear properties of Poincaré bundles that $\mathcal{L}(\mu,\nu)|_{(\mu,\nu)\in\mathfrak{n}\times\mathfrak{n}}$ determines a $\mathbb{G}_{m}$-biextension of $\underline{\mathfrak{n}}\times\underline{\mathfrak{n}}$ over $X$. Moreover, the $\mathcal{O}_{\boldsymbol{k}}$-linearality of the polarization of $B$ guarantees that $\mathcal{L}(\mu,\nu)$, up to canonical isomorphism, only depends on the image of $\mu,\nu$ in
@@
Our assumption that $D$ is odd implies that $\mathrm{Sym}_{\Phi}$ is a free $\mathbb{Z}$-module of rank $1$. There is a positive cone in $\mathrm{Sym}_{\Phi}\otimes_{\mathbb{Z}}\mathbb{R}$ uniquely determined by the condition $\mu\otimes\mu\geq 0$ for all $\mu\in\mathfrak{n}$. Thus all the line bundles $\mathcal{L}(\chi)$ are powers of the distinguished line bundle
(2.12) $\mathcal{L}_{\Phi}=\mathcal{L}(\chi_{0}),$
determined by the unique positive generator $\chi_{0}\in\mathrm{Sym}_{\Phi}$. Let $1_{(\underline{\mathfrak{n}}\times\underline{\mathfrak{n}})|_{U}}$ be the constant collection of invertible sheaves $\mathcal{O}_{\mathcal{C}_{\Phi}}$. The positivity condition means that for every $\mu\in\mathfrak{n}$, the isomorphism $\tau(\mu,\mu)$ extends (ncecessarily uniquely) to a homomorphism
$\tau(\mu,\mu):\mathcal{O}_{X}\rightarrow(c(\mu)^{\vee}\times c(\mu))^{*}(\mathcal{P}^{-1})$
and if $\mu\neq 0$, the homomorphism becomes trivial after restricting to $Z$.
There is a functor $M^{\Phi}(X,Z,U):\mathrm{DD}^{\Phi}(X,Z,U)\rightarrow\mathrm{DEG}^{\Phi}(X,Z,U)$. We briefly recall its construction. Suppose $(B,\kappa,\psi,c,c^{\vee},\tau)\in\mathrm{DD}^{\Phi}(X,Z,U)$. In particular we get a homomorphism of fppf sheaves $c^{\vee}\in\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}(\mathfrak{n},B)$ over $X$. Since
(see for example [Lan13, Proposition 3.1.5.1]), $c^{\vee}$ determines a semi-abelian scheme $(G^{\sharp})^{\vee}$ over $X$, such that there is an exact sequence of fppf sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules
(2.13) $0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathbb{G}_{m}\rightarrow(G^{\sharp})^{\vee}\xrightarrow{\pi^{\prime}}B^{\vee}\rightarrow 0.$
Similarly $c$ determines a semi-abelian scheme $G^{\sharp}$ over $X$, such that there is an exact sequence of fppf sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules
(2.14) $0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathbb{G}_{m}\rightarrow G^{\sharp}\xrightarrow{\pi}B\rightarrow 0.$
By [Lan13, Lemma 3.4.2], the condition $c=\psi\circ c^{\vee}$ guarantees that there is an $\mathcal{O}_{\boldsymbol{k}}$-linear isomorphism $\lambda^{\sharp}:G^{\sharp}\rightarrow(G^{\sharp})^{\vee}$. By [Lan13, §4.2], the datum $\tau$ in (2.11) gives us $1$-motives $M=[\underline{\mathfrak{n}}\xrightarrow{u}G^{\sharp}]$, and $M^{\vee}=[\underline{\mathfrak{n}}\xrightarrow{v}(G^{\sharp})^{\vee}]$ over $X$, where $u,v$ are morphisms of fppf sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules such that
$\pi\circ u=c^{\vee},\ \pi^{\prime}\circ v=c.$
By the proof of [BHK+20, Proposition 3.3.3], the morphism $M\rightarrow M^{\vee}$ induced by the identity map of $\mathfrak{n}$ and $\lambda^{\sharp}$ is a principal polarization of $M$ in the sense of [Del74, §10.2.11], which is compatible with the given polarization $\psi:B\rightarrow B^{\vee}$ and with the isomorphism $\mathfrak{m}\cong\mathfrak{n}^{\vee}$.
From now on we assume $R$ is normal and complete with respect to $I$, $X=\mathrm{Spec}\,R$, $Z$ is the closed subscheme $\mathrm{Spec}\,R/I$ of $X$, and $U$ is the open subscheme $X\setminus Z$ of $X$. Let us recall Mumford’s construction (see [Lan13, §4.5]). Let $^{\heartsuit}G$ be the (analytic) quotient of $G^{\sharp}$ by the image of the period map $\underline{\mathfrak{n}}\xrightarrow{u}G^{\sharp}$. Then the $\mathcal{O}_{\boldsymbol{k}}$-action $\kappa$ descends to an $\mathcal{O}_{\boldsymbol{k}}$-action $^{\heartsuit}\iota:\mathcal{O}_{\boldsymbol{k}}\rightarrow\operatorname{End}({}^{\heartsuit}G)$, and the principal polarization of $M$ gives us the principal polarization of $^{\heartsuit}G$. The positivity of $\tau$ guarantees that we get a degenerating abelian scheme $({}^{\heartsuit}G,{}^{\heartsuit}\iota,{}^{\heartsuit}\lambda)$ relative to $(X,Z,U)$.
Theorem 2.8.
Assume that $R$ is a Noetherian domain complete with respect to an ideal $I$ satisfying $\mathrm{rad}(I)=I$, and
$(X,Z,U)=(\mathrm{Spec}\,R,\mathrm{Spec}\,(R/I),\eta)$
where $\eta$ is the generic point of $\mathrm{Spec}\,R$. Then we have an equivalence of categories
(2.15) $M^{\Phi}(X,Z,U):\mathrm{DD}^{\Phi}(X,Z,U)\rightarrow\mathrm{DEG}^{\Phi}(X,Z,U)$
Moreover, this restricts to an equivalence of categories
$\mathrm{DD}^{\Phi}_{(n-1,1)}(X,Z,U)\rightarrow\mathrm{DEG}^{\Phi}_{(n-1,1)}(X,Z,U).$
Proof.
The fact that $M^{\Phi}(X,Z,U)$ is an equivalence of category is a special case of [Lan13, Theorem 5.1.1.4]. It remains to show that $\mathrm{DEG}^{\Phi}_{(n-1,1)}(X,Z,U)$ is the essential image of the functor
$\mathrm{Id}\times M^{\Phi}(X,Z,U):\mathcal{M}_{(1,0)}(X)\times\mathrm{DD}^{\Phi}_{(n-1,1)}(X,Z,U)\rightarrow\mathcal{M}_{(1,0)}(X)\times\mathrm{DEG}^{\Phi}_{(n-1,1)}(X,Z,U)$
when restricted on $\mathrm{DD}^{\Phi}_{(n-1,1)}(X,Z,U)$. This is true by [How15, Lemma 2.3.5] and [How15, Lemma 2.3.6]. ∎
2.6. Formal boundary charts
We describe the boundary more explicitly following [BHK+20, §3.3]. For $\left(A_{0},B,\ldots\right)\in\mathcal{M}_{(1,0)}(S)\times_{\mathcal{O}_{\boldsymbol{k}}}\mathcal{M}_{(n-2,0)}(S)$, the étale sheaf $\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}\left(B,A_{0}\right)$ is locally constant by [BHY15, Theorem 5.1]. For a fixed cusp label $\Phi=(\mathfrak{n},L_{\Phi})$, let $\Lambda_{\Phi}=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(L_{\Phi},\mathcal{O}_{\boldsymbol{k}})$ be the hermitian dual of $L_{\Phi}$. Let $\mathcal{A}_{\Phi}$ be the moduli space of triples $\left(A_{0},B,\varrho\right)$ over $\mathcal{O}_{\boldsymbol{k}}$-schemes $S$, where
$\left(A_{0},\ldots,B,\ldots\right)\in\mathcal{M}_{(1,0)}(S)\times_{\mathcal{O}_{\boldsymbol{k}}}\mathcal{M}_{(n-2,0)}(S),$
and
is an isomorphism of étale sheaves of hermitian $\mathcal{O}_{\boldsymbol{k}}$-modules. Then $\mathcal{A}_{\Phi}\rightarrow\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$ is smooth of relative dimension 0. Now we define $\mathcal{B}_{\Phi}$ to be the moduli space of quadruples $\left(A_{0},B,\varrho,c^{\vee}\right)$, where for an $\mathcal{O}_{\boldsymbol{k}}$-schemes $S$ we have $\left(A_{0},B,\varrho\right)\in\mathcal{A}_{\Phi}(S)$ and $c^{\vee}:\mathfrak{n}\rightarrow B$ is an $\mathcal{O}_{\boldsymbol{k}}$-linear homomorphism of group schemes over $S$. In other words,
where $\left(A_{0},B,\varrho\right)$ is the universal object over $\mathcal{A}_{\Phi}$. More explicitly, according to [BHK+20, Proposition 3.4.4], we have
(2.17) $\mathcal{B}_{\Phi}\cong E\otimes_{\mathcal{O}_{\boldsymbol{k}}}L_{\Phi},$
where $E=\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}(\mathfrak{n},A_{0})\in\mathcal{M}_{(1,0)}(S)$ and $\otimes$ is Serre’s tensor construction. The forgetful morphism $\mathcal{B}_{\Phi}\rightarrow\mathcal{A}_{\Phi}$ is smooth of relative dimension n2n-2.